Two spherical objects have a combined mass of 190 kg . The gravitational attraction between them is 8.03×10−6 N when their centers are 22.0 cm apart.

Part A

What is the mass of the heavier object?

Express your answer with the appropriate units.

nothingnothing

SubmitRequest Answer

Part B

What is the mass of the lighter object?

Express your answer with the appropriate units.

nothingnothing
Answer #1

Here,

let the masses are x and y

as the sum of the masses is 190 kg

x + y = 190 ---(1)

for the force between the masses

8.03 *10^-6 = 6.673 *10^-11 * x * y/0.22^2 ---(2)

solving 1 and 2

x = 151.6 Kg

y = 38.4 Kg

part A)

the heavier mass is 151.6 Kg

part B)

the lighter mass is 38.4 Kg